Home‎ > ‎Reports‎ > ‎Crew 108‎ > ‎

day04


Crew #108 Log Book for December 06, 2011

Photos of the day






Commander's Report


Date and time:

6 December 2011


Written by : 

Charlotte Poupon, Commander


Today we made 3 EVAs.


At the beginning of the afternoon, we welcomed the movie crew of Alicia, our mission Artist. After presenting our Hab, we went to Land Art Creek - as documented in EVA # 2 - to shoot some images. This EVA was with 6 people, after authorization by Mission Support. We took the 3 ATV and the pressurized rover.

After a security check, the commander was the last one to exit the Hab to lock it behind her.

We came back at 7:20 and started to cook dinner, as it is a cooking day.

Engineering Report

Written by:

Ashley Dale, Engineer


Date:

6 December 2011


Time:

9:00 pm

Inverter/ Batteries
Heat Tapes Operating 24/7
One Meter
Amps Where is this?
Volts Where is this?
Watts Where is this?
kWHr Where is this?
Vehicles
New Blue
Used 3/4 of tank left
Oil not checked
Tyres good
Odometer (miles) not checked

ATV 1
Gasoline used (gallons) 2/3
Oil Checked yes
Tyres good

ATV 2
Gasoline used (gallons) 1/2
Oil Checked yes
Tyres good

ATV 3
Gasoline used (gallons) 2/3
Oil Checked yes
Tyres good

Water
Trailer (inches from bottom) 0.5
Outside Potable (inches from top) 8
Inside Potable (markings) 8 inches

Temperature (deg F)
Min Don't know
Current Is there an external temperature sensor?
Max Don't know
Humidity (%)
Min Don't know
Current 10
Max Don't know

Observatory
radio telescope not used
Musk observatory not used
Hab
HVAC OK
Mars Surface Suits OK except for loose joints on 4 and 5
Computers/Network Ok except webcams not uploading
Appliances OK
GreenHab Offline
Observatory not used


Upgrades Fixed up EVA suit 2 joint (super glue)
Fixed up battery on EVA suit 4
Notes Need to repair suit 2 - loose battery

Journalist Report

Date and time: 

6 December 2011


Written by:
Ashley Dale, Engineer

Jetlag finally subsiding for us, Europeans, waking up to the smell of Usha making bread and baking blackberry muffins after a good night sleep was brilliant. That lady is a God-send. Sunrise was particularly captivating this morning – a couple of us watched it from the roof of the Hab as it washed in over the hill to our east. The crew is bonding really well. Meals together are always a group effort and a lot of fun. EVAs are now well organized, and quickly executed – we’re making sure not to explore the same regions twice, for now. We’ve gotten into a tradition of going out in the evenings to look at the stars with cups of tea at hand.

EVA 6 Report

Date and time:

6 December 2011


Written by: 

Michael LeClair


Time:

09:03 am -10:51am


Crew Members:

Poupon(Sortie Lead), LeClair (Geologist), Lingappa (Biologist)


Site Location:

Traverse 1:  Lowell Hwy to North of Ham Hillock.  Dead-reckoning nav used to locate small water-body from EVA 3.  Climbed small ridge, returned to vehicles.

 

Traverse 2:         

Stop 1.  Water-body

Water samples taken, -5 degrees C (OAT), ice thickness ~5cm.  Reversed track, crossed Hwy to NE. 

Stop 2.  Volcanic sample.  Soil samples.  Soil samples with possible bacteria; also rock sample with 0.5 to 4 cm sized concretions.   

 

Traverse 3:  


Climbed hill named it Santa Claus Hill; COMM test positive.    River bed bank at road provided ample sheety NaCl samples.

 


Transit Mode:

ATV/Foot


Objectives:

Scientific Exploration


Results:

Traverse 2:         


Stop 1.  Water samples taken, -5 degrees C (OAT), ice thickness ~5cm

Stop 2.   Volcanic and Soil samples with possible bacteria. Rock sample with 0.5 to 4 cm sized concretions. Canyon for rover exploration identified.


Traverse 3:  


COMM test positive from hill. River bed bank at road provided ample sheety NaCl samples.


EVA 7 Report

Date and time:

6 December 2011


Written by:

Michael Lotto


Time:

12:00am - 2:00pm


Crew Members:

Lotto (Sortie Lead) and Dale (Engineer)


Site Location:

Traverse 1: Ridgeline and mountains North of Hab

Traverse 2: Plains East of Hab

Transit Mode:

Foot/ATV

Objectives:

Exploration and basic geology


Results:

- Found several interesting geological samples on Traverse 1
- Lost two rock hammers along road but found both
- Found future EVA site East of Hab on Traverse 2


Lessons Learned:

EVA 8 Report

Date and time:

6 December 2011


Written by: 

Usha Lingappa


Time:

3:15pm - 5:10pm


Crew Members:

Lotto, Dale, Lingappa, LeClair, Poupon, Framis, & film
crew form Amsterdam.


Site Location:

Area with white rocks that Lotto, Poupon, and Dale found on EVA 2.

Transit Mode:

Lotto, Dale, and Lingappa traveled by ATV.
Poupon, LeClair, and Framis traveled by HabCar (pressurized rover).



Objectives:

First shoot with Alicia's film crew.


Results:

Film crew obtained several shots of the whole crew climbing, walking in the sunset, and using Alicia's moon compasses.


Lessons Learned: